Section 8.1: Problem 3 Previous Problem Problem List Next Problem (1 point) Take the conservative equation X + x2 - 64 = 0.
Answer #1

xa x²64 = 0 » X=64-22 9 = 64-27 For caulical points on x axis. (y=0) put y: 0, => 64-? = 0 = 2 = t8 : x = [78] & 1-8] : ThUND Ś JE TO 17 1-2x of At, 2=8. to 17 1-16 od. Let eigenvalue = a. (a)x( a) + 16 = 0 => 72--16 : => fi.4 So, this is centreo)